I need help with the following distributive problem:
6n-3=8n+12

Answer by checkley71(8403)

6N-3=8N+12 OR 6N-8N=12+3 OR -2N=15 OR N=15/-2 05 N=-7.5
PROOF 6*-7.5-3=8*-7.5+12 OR -45-3=-60+12 OR -48=-48